# DeNOTS: Stable Deep Neural ODEs for Time Series

### v4

@@ 1 Introduction
Neural Controlled Differential Equations (CDEs) Kidger et al. (2020) provide a natural way to process irregular time series. CDEs are Ordinary Differential Equations (ODEs), where the derivative depends on an external input signal. Neural CDEs utilize a Neural Network (NN) as the ODE’s dynamics function. It is well-known that increasing NN depth (number of layers) leads to higher expressiveness Gripenberg (2003); Lu et al. (2017); Yarotsky (2017), i.e., widens the class of functions the NN may represent Gühring et al. (2020). According to the original Neural ODE paper Chen et al. (2018), the natural analogue of NN depth is the number of function evaluations (NFE). Naturally, we hypothesise that a larger NFE results in a better Neural CDE model.
@@
NFE is primarily controlled by the solver tolerance, which defines the acceptable error level during numerical integration Dormand & Prince (1980). Lowering this tolerance increases the required number of integration steps and NFE. However, prior work mostly sidesteps this topic: the expressiveness gains from higher precision are often minimal in practice. The problem here is that, as we will show, boosting expressiveness on a fixed integration interval necessitates larger $l_{2}$ weight norms, which harms training stability. Instead, we propose scaling the integration time. This method improves expressiveness while reducing the required weight norms. We refer to it as Scaled Neural CDE (SNCDE).
@@
Upon investigating the proposed time-scaling procedure, we found that longer integration intervals can introduce uncontrollable trajectory growth, which must also be addressed. A very intuitive approach to constraining the trajectory is adding Negative Feedback (NF). Prior work implemented it by subtracting the current hidden state from the dynamics function De Brouwer et al. (2019). However, as we demonstrate, such a technique causes ”forgetfulness“: the influence of older states constantly decays, and the model cannot retain important knowledge throughout the sequence. This effect is akin to the one experienced by classic Recurrent Neural Networks (RNN), remedied by Long Short-Term Memory (LSTM) Hochreiter & Schmidhuber (1997) and Gated Recurrent Units (GRU) Chung et al. (2014). We demonstrate that our novel NF does not suffer from ”forgetfulness“.
@@
Time series models possess another crucial property: we work with discrete observations instead of the true continuous process. The models cannot analyse the system during the resulting inter-observation gaps, accumulating epistemic uncertainty in the final representation Blasco et al. (2024). Estimating such uncertainty has been a topic of interest for decades Golubev & Krymova (2013). Without additional modifications, it piles up with time, and the theoretical variance of the final prediction becomes proportional to sequence length. High variance makes the model vulnerable to various adversarial samples Cohen et al. (2019). In our case, the model is vulnerable to attacks that increase the intervals between input sequence elements by randomly dropping observations.
@@
  • The antisynchronous NF mechanism (updates and NF are activated in anti-phase), which stabilises hidden trajectories on larger time scales but keeps the model flexible in practice. All alternatives have crucial flaws, as summarized in Table 1: non-stable dynamics are challenging to train, and synchronous NF tends to ”forget“ important information.
  • Theoretical analysis of the effect of discretisation on our model, certifying that epistemic uncertainty does not accumulate in our prediction. We also provide an analytical bound for our case of cubic spline interpolation.
@@
SNCDE Vector Field (Th. 4.4/Fig. 7) (Th. 4.5/Tab. 7) (Th. 4.9/Tab. 6)
@@ 2 Method
The output of our backbone is $\mathbf{h}\triangleq\mathbf{h}(T)$. We use a modified GRU cell Chung et al. (2014) for $\mathbf{g}_{\theta}$, and, following Kidger et al. (2020), interpolate the input data using natural cubic splines.
@@
    Interpolate the multivariate features $\{\mathbf{x}_{k}\}_{k}$ to get $\hat{\mathbf{x}}(t),t\in[0,T]$ using cubic splines Kidger et al. (2020).
@@ 3 Time Scaling
As mentioned in the introduction, we argue that scaling time benefits expressiveness. By expressiveness, we roughly mean the broadness of the class of functions that our network can represent. Formally, our model $F$ maps the continuous input data trajectory $\hat{\mathbf{x}}(\cdot):[0,T]\rightarrow\mathbb{R}^{u}$ to a hidden state $\mathbf{h}\triangleq\mathbf{h}(T)\in\mathbb{R}^{v}$, i.e. $F:C([0,T],\mathbb{R}^{u})\rightarrow\mathbb{R}^{v}$. We describe the class of functions $F$ in terms of Lipschitz constants, where $F$ is $L_{F}$-Lipschitz, if:
$\|F(\hat{\mathbf{x}}_{1})-F(\hat{\mathbf{x}}_{2})\|_{2}\leq\rho(\hat{\mathbf{x}}_{1},\hat{\mathbf{x}}_{2})L_{F},\;\;\;\textrm{where}\;\rho(\hat{\mathbf{x}}_{1},\hat{\mathbf{x}}_{2})\triangleq\int_{0}^{T}\|\hat{\mathbf{x}}_{1}(t)-\hat{\mathbf{x}}_{2}(t)\|_{2}dt.$ (2)
Note that in this section, for simplicity, we assume that $\mathbf{h}(0)=\mathbf{0}_{v}$. To broaden the class of functions the model may represent, we must increase $L_{F}$. We can explicitly bound $L_{F}$, using $T$ and the Lipschitz constants of $\mathbf{g}_{\theta}$:
Theorem 3.1.
Suppose, that the function $\mathbf{g}_{\theta}$ is Lipschitz w.r.t. $\hat{\mathbf{x}},\mathbf{h}$ with constants $M_{x},M_{h}$ respectively. Then, $F$ is $L_{F}$-Lipschitz with $L_{F}=\frac{M_{x}}{M_{h}}(e^{M_{h}T}-1)$.
Consequently, one way to boost expressivity is to increase $M_{h}$, corresponding to the weights’ norms. Larger weight norms induce saturation for bounded activations such as tanh or sigmoid Ven & Lederer (2021) or ”dying“ for unbounded activations such as ReLU Lu et al. (2019), destabilising training in both cases. Alternatively, scaling $T$ makes the model exponentially more expressive without increasing the magnitude of weights. We refer to Neural CDEs with time scaling as ”Scaled Neural CDEs“ (SNCDEs). Formally, time scaling means we set $t_{k}\leftarrow\frac{D}{M}t_{k}$, where $D>0$ is a hyperparameter and $M>0$ is a dataset-specific normalising constant, introduced to make values of $D$ comparable across benchmarks.
@@ 4 Negative Feedback
In practice, the right-hand side of 3 is a slightly modified version of the classic GRU architecture Chung et al. (2014). Specifically, we work with the last step in the GRU block, which is:
@@
The GRU-ODE version, proposed in De Brouwer et al. (2019), which subtracts $\mathbf{h}$ from (5). Here, both the update term $\mathbf{n}$ and the NF term $\mathbf{h}$ are activated simultaneously, with $\mathbf{z}$ affecting only the overall scale of the derivative, so we refer to this version as Sync-NF. The $l_{2}$ norms of the weights are smaller on longer integration intervals (as we concluded in the previous section), so Assumption 4.2 always stands for Sync-NF. The Sync-NF version without time scaling ($D=1$) is equivalent to GRU-ODE [^2] From here on: GRU-ODE refers to the corresponding NCDE without time scaling ($D=1$). De Brouwer et al. (2019). As its authors prove, the corresponding trajectory is constrained to [1,+1][-1,+1]. Moreover, it does not regulate the scale of NF relative to the update, implying that, intuitively, the model may ”forget“ prior knowledge (see Section 4.3).
@@
We pass $\mathbf{h}$ instead of $\mathbf{h}$ to the GRU layer. In our version, the NF and update terms are activated separately, with $\mathbf{z}$ regulating the scale of NF relative to the update, so we refer to this version as Anti-NF. The validity of (4) for Anti-NF depends on the value of $a$. Consequently, our model can disable the ”forgetting“ effect by increasing $a$, which is learnable and even adaptive, since it corresponds to $\mathbf{z}$ from the GRU (5).
@@ 4.1 Stability
A crucial issue to consider when working with ODEs is stability analysis Boyce et al. (1969); Oh et al. (2024). On longer time frames, unstable ODEs may have unbounded trajectories even with reasonable weight norms, which is detrimental to training due to saturation Ioffe & Szegedy (2015) or ”dying“ Lu et al. (2019) of activations. The equation (1) depends on an input signal $\hat{\mathbf{x}}(t)$, so we approach stability analysis in terms of control systems theory Sontag & Wang (1995):
@@ 4.2 Error Bounds
The approximation $\hat{\mathbf{x}}(t)$ of the true unknown $\mathbf{x}(t)$ from discrete observations is a source of errors. As we will show below, NF significantly helps reduce the effect of these errors. The form (3) allows us to measure how this discrepancy affects the quality of our final embedding $\mathbf{h}(T)$.
We assume that $\mathbf{x}(t)$ is a realisation of a Gaussian process (GP) Kolmogorov (1941); Stein (2012), dense in the space of continuous functions. Cubic spline interpolation used in our algorithm and in Kidger et al. (2020) is a limit case of GP regression Golubev & Krymova (2013).
Let $\mathbf{x}(t)$ be a stationary zero-mean $u$-dimensional GP with the covariance function $\mathbf{K}(\cdot,\cdot)=\{K_{i}(\cdot,\cdot)\}_{i}$ and independent components:
$\mathbb{E}\mathbf{x}(t)=\mathbf{0};\;\;\mathbb{E}\left[\mathbf{x}(t_{1})\odot\mathbf{x}(t_{2})\right]=\mathbf{K}(t_{1},t_{2})=\mathbf{K}(t_{2}-t_{1});\;\;\mathbb{E}\left[x_{i}(t)\cdot x_{j}(t)\right]=0\textrm{ for }i\neq j.$ (7)
Here, the symbol $\odot$ denotes component-wise multiplication, and lower indexing $x_{i}$ denotes $i$-th component of a vector. Suppose that we also know the values of $\mathbf{x}$ at timestamps $0=t_{1}<t_{2}<\ldots<t_{n}=T\in\mathbb{R}$. Let $\hat{\mathbf{x}}(t)$ be the corresponding maximum a posteriori predictor Kolmogorov (1941). We introduce two types of uniform error bounds:
$\sigma_{\mathrm{PW}}^{2}\triangleq\mathbb{E}\left\lVert\hat{\mathbf{x}}(t)-\mathbf{x}(t)\right\rVert_{2}^{2}\leq\bar{\sigma}_{\mathrm{PW}}^{2},$ $\sigma_{\mathrm{INT}}^{2}\triangleq\int_{t_{k}}^{t_{k+1}}\mathbb{E}\left\lVert\hat{\mathbf{x}}(t)-\mathbf{x}(t)\right\rVert_{2}^{2}dt\leq\bar{\sigma}_{\mathrm{INT}}^{2},$ (8)
where $\bar{\sigma}_{\mathrm{INT}},\bar{\sigma}_{\mathrm{PW}}$ are the pointwise and the interval error bounds, constants independent of $t$ and $k$. The pointwise expression is more convenient; however, the interval one is often tighter. We will now estimate how much of the variance from (8) accumulates in the final hidden state:
@@
Robustness (Gaussian Process standpoint). Along with (3) under Assumptions 4.14.2, consider analogous ODE dynamics for a realization of the actual GP $\mathbf{x}(t)$ instead of its MAP approximation $\hat{\mathbf{x}}(t)$, with the dynamics $\frac{d\mathbf{h}^{*}}{dt}=a\mathbf{f}_{\theta}(\mathbf{x}(t),\mathbf{h}^{*}(t))-b\mathbf{h}^{*}(t)$, and starting conditions $\mathbf{h}^{*}(0)=\mathbf{h}(0)=\mathbf{h}_{0}$. Then, depending on whether we are working with the point-wise or the interval errors (8), we can write different bounds on the variance, accumulating in $\mathbf{h}$.
Pointwise bound: $\mathbb{E}\lVert\mathbf{h}(t_{n})-\mathbf{h}^{*}(t_{n})\rVert_{2}^{2}\leq\left(\frac{aL_{x}}{b-aL_{h}}\right)^{2}\bar{\sigma}_{\mathrm{PW}}^{2}.$ (9)
Interval bound: $\frac{1}{n}\int_{0}^{T}\mathbb{E}\left\lVert\mathbf{h}(t)-\mathbf{h}^{*}(t)\right\rVert_{2}^{2}dt\leq\left(\frac{aL_{x}}{b-aL_{h}}\right)^{2}\bar{\sigma}^{2}_{\mathrm{INT}}.$ (10)
The errors (9) and (10) are independent of the number of observations $n$. Consequently, uncertainty does not accumulate in our final hidden state.
We empirically validate the benefits of Theorem 4.5 with drop attacks (dropping a fraction of input tokens and replacing them with NaNs) on the Pendulum dataset. According to Table 7 (first row), all SNCDEs are robust to drop attacks, with NF versions outperforming No NF. The Sync-NF model is slightly better than Anti-NF, likely due to a larger $b-aL_{h}$, but that difference is insignificant compared to other baselines. Additional analysis of tightness of (9), (10) is in Appendix B.4.7.
4.2.1 Error bounds for specific covariance functions
Most prior theoretical results on GP interpolation are devised for the case of uniform sequences. However, they can be extended to irregular data, assuming that the resulting error increases in response to stretching intervals $\delta_{k}\triangleq t_{k+1}-t_{k}$. Then, the upper and lower bounds for uniform sequences are also valid for irregular sequences, given that instead of the regular step $\delta\equiv\delta_{k}$ we take the largest interval $\delta_{\max}\triangleq\max_{k}(\{\delta_{k}\}_{k})$, or the smallest one $\delta_{\min}\triangleq\min_{k}(\{\delta_{k}\}_{k})$ respectively.
Additionally, prior work Kolmogorov (1941) assumes that the considered sequences are infinite. However, the authors of Zaytsev & Burnaev (2017) demonstrated that the results for infinite sequences empirically hold for finite ones. Indeed, a typical kernel decays exponentially, so distant observations negligibly impact the resulting error.
Given the reasoning above, the following assumption does not limit the applicability of our conclusions, while allowing us to devise GP error bounds:
Assumption 4.6.
The considered sequence $S$ is infinite, and the covariance function is such that increasing interval sizes increases the interval error from 8.

### v5

@@ 1 Introduction
Neural Controlled Differential Equations (CDEs) kidger2020neural provide a natural way to process irregular time series. CDEs are Ordinary Differential Equations (ODEs), where the derivative depends on an external input signal. Neural CDEs utilize a Neural Network (NN) as the ODE’s dynamics function. It is well-known that increasing NN depth (number of layers) leads to higher expressiveness gripenberg2003approximation; lu2017expressive; yarotsky2017error, i.e., widens the class of functions the NN may represent guhring2020expressivity. According to the original Neural ODE paper chen2018neural, the natural analogue of NN depth is the number of function evaluations (NFE). Naturally, we hypothesise that a larger NFE results in a better Neural CDE model.
@@
NFE is primarily controlled by the solver tolerance, which defines the acceptable error level during numerical integration dormand1980family. Lowering this tolerance increases the required number of integration steps and NFE. However, prior work mostly sidesteps this topic: the expressiveness gains from higher precision are often minimal in practice. The problem here is that, as we will show, boosting expressiveness on a fixed integration interval necessitates larger $l_{2}$ weight norms, which harms training stability. Instead, we propose scaling the integration time. This method improves expressiveness while reducing the required weight norms. We refer to it as Scaled Neural CDE (SNCDE).
@@
Upon investigating the proposed time-scaling procedure, we found that longer integration intervals can introduce uncontrollable trajectory growth, which must also be addressed. A very intuitive approach to constraining the trajectory is adding Negative Feedback (NF). Prior work implemented it by subtracting the current hidden state from the dynamics function de2019gru. However, as we demonstrate, such a technique causes ”forgetfulness“: the influence of older states constantly decays, and the model cannot retain important knowledge throughout the sequence. This effect is akin to the one experienced by classic Recurrent Neural Networks (RNN), remedied by Long Short-Term Memory (LSTM) hochreiter1997long and Gated Recurrent Units (GRU) chung2014empirical. We demonstrate that our novel NF does not suffer from ”forgetfulness“.
@@
Neural differential equation-based models possess another crucial property: we cannot calculate the solution perfectly, forcing us to approximate the true continuous process via discrete steps. We have to discretise the time series, so the models cannot analyse the system during the resulting inter-observation gaps, accumulating epistemic uncertainty in the final representation BLASCO2024127339. Estimating such uncertainty has been a topic of interest for decades golubev2013interpolation. Differential equation solvers also introduce discretisation errors, since they simulate a difference equation instead of the continuous differential one hairer1993solving. Without additional modifications, the error piles up with time, and the theoretical variance of the final prediction becomes proportional to sequence length.
@@
  • The antisynchronous NF mechanism (updates and NF are activated in anti-phase), which stabilises hidden trajectories on larger time scales but keeps the model flexible in practice. All alternatives have crucial flaws, as summarized in Table 1: non-stable dynamics are challenging to train, and synchronous NF is to restricitve, and tends to ”forget“ important information.
  • Theoretical analysis of the effect of discretisation on our model, certifying that epistemic uncertainty does not accumulate in our prediction. We also provide a tight analytical bound for the interpolation error of natural cubic splines.
@@
SNCDE Vector Field (Th. 4.4/Fig. 7) (Th. 4.5/Tab. 7) (Th. 4.8/Tab. 6)
@@ 2 Method
The output of our backbone is $\mathbf{h}\triangleq\mathbf{h}(T)$. We use a modified GRU cell chung2014empirical for $\mathbf{g}_{\theta}$, and, following kidger2020neural, interpolate the input data using natural cubic splines.
@@
    Interpolate the multivariate features $\{\mathbf{x}_{k}\}_{k}$ to get $\hat{\mathbf{x}}(t),t\in[0,T]$ using cubic splines kidger2020neural.
@@ 3 Time Scaling
As mentioned in the introduction, we argue that scaling time benefits expressiveness. By expressiveness, we mean the broadness of the class of functions that our network can represent. Now, we introduce several definitions.
Differentially definable mappings from trajectories $\mathcal{F}$.
We say that $F_{\mathbf{g}}\in\mathcal{F}$, if there exists an ODE with the vector field $\mathbf{g}(\mathbf{x}(t),\mathbf{h}(t)):\mathbb{R}^{u}\times\mathbb{R}^{v}\rightarrow\mathbb{R}^{v}$, with the starting condition $\mathbf{h}(0)=\mathbf{h}_{0}$ and the solution $\mathbf{h}(t):\mathbb{R}\rightarrow\mathbb{R}^{v}$: $\frac{d}{dt}\mathbf{h}(t)=\mathbf{g}(\mathbf{h}(t),\mathbf{x}(t)),$ such that $F_{\mathbf{g}}(\mathbf{x}(\cdot),\mathbf{h}_{0};t)=\mathbf{h}(t)$. Not all functions on trajectories can be represented this way.
Start-corrected differentially-definable mappings,
i.e. the elements of $\mathcal{F}$, minus their initial state $\mathbf{h}_{0}$: $\mathring{\mathcal{F}}=\{\mathring{F}_{\mathbf{g}}(\mathbf{x}(\cdot),\mathbf{h}_{0};t)=F_{\mathbf{g}}(\mathbf{x}(\cdot),\mathbf{h}_{0};t)-\mathbf{h}_{0}|F_{\mathbf{g}}\in\mathcal{F}\}.$ Since we set $\mathbf{h}_{0}=0$ in most practical cases, the distinction is rather technical. This is mostly to account for Lipschitzness w.r.t. $\mathbf{h}$, as we will see below.
Lipschitz constraints.
We introduce two ways to constrain this class with Lipschitz constants: with $\mathbf{g}$ being $M_{x},M_{h}$-Lipschitz
$\mathring{\mathcal{F}}_{g}(M_{x},M_{h})=\{\mathring{F}_{\mathbf{g}}\in\mathring{\mathcal{F}}|\mathbf{g}-M_{x},M_{h}\text{-lipschitz w.r.t. }\mathbf{x},\mathbf{h}\},$
and with $\mathring{F}_{\mathbf{g}}(\ldots;t)$ being $L_{x}(t),L_{h}(t)$-lipschitz, where $L_{x}(t),L_{h}(t):\mathbb{R}\rightarrow\mathbb{R}_{+}$:
$\mathring{\mathcal{F}}_{F}(L_{x}(\cdot),L_{h}(\cdot))=\{\mathring{F}_{\mathbf{g}}\in\mathring{\mathcal{F}}|\mathring{F}_{\mathbf{g}}-L_{x}(t),L_{h}(t)\text{-lipschitz w.r.t. }\mathbf{x},\mathbf{h}_{0}\}.$
Note, that $\mathring{\mathcal{F}}_{g}(M_{x},M_{h})$ represents all mappings that our ODE can learn, given a Lipschitz-constrained Neural Network $\mathbf{g}$, while $\mathring{\mathcal{F}}_{F}(L_{x}(\cdot),L_{h}(\cdot))$ represents all the possible Lipschitz-mappings that we may want to learn.
Equipped with the above definitions, we can formulate the following:
Theorem 3.1 (Expressiveness).
The classes $\mathring{\mathcal{F}}_{g}(M_{x},M_{h})$ and $\mathring{\mathcal{F}}_{F}(L_{x}(\cdot),L_{h}(\cdot))$ are equal, given the following relations between their arguments:
$L_{x}(t)=M_{x}\sqrt{\frac{1}{2M_{h}}\left(e^{2M_{h}t}-1\right)};L_{h}(t)=e^{M_{h}t}-1.$
As a corollary, given an $M_{x},M_{h}$-Lipschitz NN $\mathbf{g}$, we can represent all the $L_{x},L_{h}$-Lipschitz (differentially-definable, start-corrected) $F$, and only them. Consequently, one way to boost expressivity is to increase $M_{h}$, corresponding to the weights’ norms. Larger weight norms induce saturation for bounded activations such as tanh or sigmoid ven2021regularization or ”dying“ for unbounded activations such as ReLU lu2019dying, destabilising training in both cases. Alternatively, scaling $T$ makes the model exponentially more expressive without increasing the magnitude of weights. We refer to Neural CDEs with time scaling as ”Scaled Neural CDEs“ (SNCDEs). Formally, time scaling means we set $t_{k}\leftarrow\frac{D}{M}t_{k}$, where $D>0$ is a hyperparameter and $M>0$ is a dataset-specific normalising constant, introduced to make values of $D$ comparable across benchmarks.
@@ 4 Negative Feedback
In practice, the right-hand side of 2 is a slightly modified version of the classic GRU architecture chung2014empirical. Specifically, we work with the last step in the GRU block, which is:
@@
The GRU-ODE version, proposed in de2019gru, which subtracts $\mathbf{h}$ from (4). Here, both the update term $\mathbf{n}$ and the NF term $\mathbf{h}$ are activated simultaneously, with $\mathbf{z}$ affecting only the overall scale of the derivative, so we refer to this version as Sync-NF. The $l_{2}$ norms of the weights are smaller on longer integration intervals (as we concluded in the previous section), so Assumption 4.2 always stands for Sync-NF. The Sync-NF version without time scaling ($D=1$) is equivalent to GRU-ODE [^2] From here on: GRU-ODE refers to the corresponding NCDE without time scaling ($D=1$). de2019gru. As its authors prove, the corresponding trajectory is constrained to [1,+1][-1,+1]. Moreover, it does not regulate the scale of NF relative to the update, implying that, intuitively, the model may ”forget“ prior knowledge (see Section 4.3).
@@
We pass $\mathbf{h}$ instead of $\mathbf{h}$ to the GRU layer. In our version, the NF and update terms are activated separately, with $\mathbf{z}$ regulating the scale of NF relative to the update, so we refer to this version as Anti-NF. The validity of (3) for Anti-NF depends on the value of $a$. Consequently, our model can disable the ”forgetting“ effect by increasing $a$, which is learnable and even adaptive, since it corresponds to $\mathbf{z}$ from the GRU (4). Setting $a=1,b=0$ lifts the NF constraint altogether, turning (2) into a generic Neural CDE: $\frac{d\mathbf{h}}{dt}=\mathbf{f}_{\theta}$.
@@ 4.1 Stability
A crucial issue to consider when working with ODEs is stability analysis boyce1969elementary; oh2024stable. On longer time frames, unstable ODEs may have unbounded trajectories even with reasonable weight norms, which is detrimental to training due to saturation ioffe2015batch or ”dying“ lu2019dying of activations. The equation (1) depends on an input signal $\hat{\mathbf{x}}(t)$, so we approach stability analysis in terms of control systems theory sontag1995characterizations:
@@ 4.2 Error Bounds
Until now, we have assumed that the Neural Differential Equation (1) which our method models can be solved perfectly. Unfortunately, due to physical restrictions of our computational devices, obtaining the perfect solution is impossible. We are forced to discretise the continuous processes that we work with, which introduces certain errors into our calculations:
  • Discretisation of the input process $\hat{\mathbf{x}}(t)$ introduces the interpolation error. The interpolation error is a property of the interpolation routine.
  • Discretisation of the numerical integration process introduces the integration error. The integration error is a property of the differential equation solver.
It is a well-known result hairer1993solving, that the global numerical integration error is exponential w.r.t. time. In a sense, we observe precisely that in Figure 6, where the hidden norms explode for non-stable vector fields. The exponential accumulation of error significantly hampers the time-scaling routine that we propose, so training such models becomes close to impossible. Fortunately, we were able to prove that our stable Anti-NF and Sync-NF models, adhering to Assumption 4.2, do not accumulate the numerical error in their final hidden state.
To devise the theoretical results of this section, we need to introduce the ”perfect“ Neural CDE, calculated without any errors:
$\mathbf{h}^{*}(0)=\mathbf{h}_{0};\quad\frac{d\mathbf{h}^{*}(t)}{dt}=a\mathbf{f}_{\theta}(\mathbf{x}(t),\mathbf{h}^{*}(t))-b\mathbf{h}^{*}(t).$ (6)
Then, both the integration and the interpolation errors may be modelled via the following ”imperfect“ ODE, often referred to as the ”Modified Vector Field“ reich1999backward:
$\mathbf{h}(0)=\mathbf{h}_{0};\quad\frac{d\mathbf{h}(t)}{dt}=a\mathbf{f}_{\theta}(\mathbf{x}(t),\mathbf{h}(t))-b\mathbf{h}(t)+\xi(t),$ (7)
where $\xi(t)$ is a random vector, which represents the momentary deviation of our numerical ODE from the true one. For further theoretical considerations, we introduce two types of bounds on $\xi(t)$:
$\text{Pointwise: }\mathbb{E}\left\lVert\xi(t)\right\rVert_{2}^{2}\leq\xi_{\mathrm{PW}}^{2},\forall 0\leq t\leq T;$ $\text{Interval: }\frac{1}{T}\int_{0}^{T}\mathbb{E}\left\lVert\xi(t)\right\rVert_{2}^{2}dt\leq\xi_{\mathrm{INT}}^{2}.$ (8)
For each of the introduced bounds, we can devise an estimate for the error in the final hidden state:
@@
Error Bounds (Robustness). Consider the differential equations (6),(7), under Assumptions 4.14.2 Then, depending on whether we are working with the pointwise or the interval errors (8), we can write different bounds on the variance, accumulating in $\mathbf{h}$.
Pointwise bound: $\mathbb{E}\lVert\mathbf{h}(T)-\mathbf{h}^{*}(T)\rVert_{2}^{2}\leq\left(\frac{a}{b-aL_{h}}\right)^{2}\xi_{\mathrm{PW}}^{2}.$ (9)
Interval bound: $\frac{1}{T}\int_{0}^{T}\mathbb{E}\left\lVert\mathbf{h}(t)-\mathbf{h}^{*}(t)\right\rVert_{2}^{2}dt\leq\left(\frac{a}{b-aL_{h}}\right)^{2}\xi^{2}_{\mathrm{INT}}.$ (10)
The errors (9) and (10) are independent of the total integration time $T$. Consequently, the error does not accumulate in our final hidden state. Additional analysis of tightness of (9), (10) is in Appendix B.4.1.
Numerical integration error is directly connected to our momentary deviation $\xi(t)$, and, as we mentioned above, the benefits of NF for solution stability are already validated by Figure 6. However, the relation of interpolation error to $\xi$ is less straightforward, we discuss it in the following section.
4.2.1 Interpolation error
Say, $\hat{\mathbf{x}}(t)$ is the cubic-spline interpolation of the input time series. For analytic purposes, we also introduce the true (unobserved) value of the input process, $\mathbf{x}(t)$. Similar to (8), we consider two types of interpolation errors:
$\left\lVert\hat{\mathbf{x}}(t)-\mathbf{x}(t)\right\rVert_{2}^{2}\leq\sigma_{\mathrm{PW}}^{2},$ $\int_{t_{k}}^{t_{k+1}}\left\lVert\hat{\mathbf{x}}(t)-\mathbf{x}(t)\right\rVert_{2}^{2}dt\leq\sigma_{\mathrm{INT}}^{2}.$ (11)
Given the absence of other errors, the momentary deviation $\xi(t)$ is connected to $\left\lVert\mathbf{x}-\hat{\mathbf{x}}\right\rVert$ via the $L_{x}$ Lipschitz coefficient, with $\xi_{\mathrm{PW/INT}}=L_{x}\sigma_{\mathrm{PW/INT}}$, so now we estimate the $\sigma$-errors. The bound for the pointwise error of cubic spline interpolation is a well-known result de1978practical:
$\|\mathbf{x}(t)-\hat{\mathbf{x}}(t)\|\leq\frac{5}{384}\delta_{\max}^{4}\|\mathbf{x}^{(4)}(t)\|,$ (12)
where $\mathbf{x}^{(4)}$ is the fourth derivative of the input trajectory, and $\delta_{\max}$ is the maximum step between observations. Assuming that $\|\mathbf{x}^{(4)}\|$ is bounded, we can plug (12) directly into (8) (remebering to include $L_{x}$).
However, the bound (12) is not tight due to the nature of cubic splines. The bound may correctly approximate the error in the middle of inter-observation gaps, but the error decreases to zero as we move towards each observation. For a more precise point of view, we can estimate the interval error from (11). We base our interval error estimation on Gaussian Process (GP) theory: the authors of golubev2013interpolation proved that, under certain conditions, cubic splines can be seen as a limit case of Gaussian Process regression. Using an approach, similar to that of zaytsev2017minimax, we were able to devise a tight bound for such a Gaussian Process limit, which we provide in the following lemma: